Q#2 (30) An aircraft flying at altitude where the density is 0.002377 slugs/ft is directed by the air traffic control to initiate a 2g constant speed level turn. The speed of the aircraft is 225 mph. (a) What is the radius of turn? (b) What is the rate of turn? Really an aerospace question.
Answer #1

Given: Load factor (LF) = 2

density of air = 0.002377 slugs/ft^3 = 0.002377*515.378818 kg/m^3 = 1.22505545039 kg/m^3

Speed of aircraft = 225mph = 362.102 kph = 100.5838889 m/s

Now,

L = sqrt{(mg)^2 + Big(mrac{v^2}{R}Big)^2} = 2mg (given LF = 2)

Therefore,

V) 3

2 3 which gives R = 343.768899295 m

Rate of turn 109 1 tan(θ) dR dt where V is the speed of aircraft in Knots and theta is the bank angle

for 2G turn, theta = 60 degrees

V = 225 mph = 195.52 knots

Rate of turn = 9.66482933233 degrees per second